At E °C, 30mL of coffee is forgotten in a thermos glass. If fresh coffee at F °C is added to the cooled coffee, the temperature of the coffee rises to G °C under room conditions. Calculate the amount of coffee added to the thermos. A two-digit number ending in 0 or 5, where E < G < F)


1
Expert's answer
2021-12-27T07:54:36-0500

"c_1V_1(G-E)=c_2V_2(F-G),"

"c_1=c_2,"

"V_2=V_1\cdot \frac{G-E}{F-G}."